OpenStudy (anonymous):

solve for d: S=a+(n-1)d

OpenStudy (anonymous):

subtract \(a\) and divide by \((n-1)\):\[S=a+(n-1)d\]\[S-a=(n-1)d\]\[{{S-a}\over{(n-1)}}=d\]\[\large d={{S-a}\over{(n-1)}}\]